Manuel's quiz scores in history are 76, 81, and 77. What score must he get on his fourth quiz to have an average of at least 80?
Answer by wrodriguez50(1)

Given: 76, 81, 77
Solve for Avg >= 80
The variable "x" is Manuel's score on the fourth quiz
(76 + 81 + 77 + x)/4 >= 80
Multiply both sides by 4 to get rid of the fraction
76 + 81 + 77 + x >= 320
Combine like terms
234 + x >= 320
Subtract both sides by 234 to get "x" alone
234 - 234 + x >= 320 - 234
X >= 86
Therefore, Manuel needs to score 86 or better on his next quiz to have an average of at least 80.